The sum of the ages of a brother and sister is 26. If four times the brother's age is subtracted from three times the sister's age, the difference is 8. What are the ages of the brother and sister, respectively?
Answer by mananth(16946)

The sum of the ages of a brother and sister is 26. If four times the brother's age is subtracted from three times the sister's age, the difference is 8. What are the ages of the brother and sister, respectively?
...
x------------- brother
y= ------------sister
x+y =26.....................1
3y-4x=8.....................2
multiply (1) by 4
we get
4x+4y = 104
add this to (2)
7y= 112
/7
y = 16 sisters age
26-16 = 10 brothers age
..
m.ananth@hotmail.ca